# Adaptive Multi-Scale Decomposition Framework for Time Series Forecasting

### v1

@@ 4.3 Adaptive Multi-predictor Synthesis Block
where $k$ is the number of dominant temporal patterns, $\psi\in\mathbb{N}(0,1)$ is standard Gaussian noise, and $\mathbf{W}_{\text{noise}}\in\mathbb{R}^{m\times m}$ is a learnable weight controlling the noisy values. The sum of the selector weights $\mathbf{S}$ for each channel is 1.
@@
TP-Projection takes the embedding $\mathbf{v}$ from the output $\mathbf{V}$ of DDI as input. It utilizes a two-layer feedforward network as a predictor. These predictions are then multiplied by the selector weights $\mathbf{S}$ and summed to yield the prediction result $\mathbf{\hat{y}}$ for one channel. The final results $\mathbf{\hat{Y}}$ are composed of the outputs $\mathbf{\hat{y}}$ from each channel:
@@
Compared to sparse MoE [40], we adopt dense MoE in TP-Projection with three considerations. Firstly, each temporal pattern contributes to the prediction result. Secondly, we take the k-th largest value instead of the k-th largest value and avoid the sorting operation by employing the divide-and-conquer method, thus reducing the time complexity from $O(nlog(n))$ to $O(n)$. Thirdly, this approach can help mitigate issues such as load unbalancing and embedding omissions that are prevalent in sparse MoE architectures. Our TopK method can be formalized with:
@@ 4.4 Loss Function
where $\|\mathbf{\Theta}\|_{2}$ is the L2-norm of all model parameters, $\lambda_{1}$ and $\lambda_{2}$ are hyper-parameters.
5 Experiments
5.1 Main Results
We thoroughly evaluate the proposed model in various time series forecasting tasks and confirm the generality of the proposed framework. Additionally, we offer insights into the effectiveness of integrating MoE components and the low computational complexity.
@@ 5.1 Main Results
We conduct experiments on seven real-world datasets, including Weather, ETT (ETTh1, ETTh2, ETTm1, ETTm2), ECL, Exchange, Traffic and Solar Energy for long-term forecasting and PEMS (PEMS03, PEMS04, PEMS07, PEMS08) for short-term forecasting. A detailed description of each dataset is provided in Sec. C.1.
@@
We carefully select some representative models to serve as baselines in the field of time series forecasting, including (1) MLP-based models: TiDE [12], MTS-Mixers [16], and DLinear [13]; (2) Transformer-based models: PatchTST [6], Crossformer [9], and FEDformer [8]; (3) CNN-based models: TimesNet [21], and MICN [20]. See Sec. C.2 for a detailed description of the baseline.
@@
To ensure fair comparisons, for long-term forecasting, we re-run all the baselines with different input lengths $L$ and choose the best results to avoid underestimating the baselines and provide a fairer comparison. For short-term forecasting, we choose an input length of 96. To trade off the memory footprint and accuracy, the number of predictors $n$ is set to 8. We select two common metrics in time series forecasting: Mean Absolute Error (MAE) and Mean Squared Error (MSE). All experiments are conducted using PyTorch on an NVIDIA V100 32GB GPU and repeat five times for consistency. See Sec. C.4 for detailed parameter settings.
Results.
Comprehensive forecasting results are shown in Tab. 1 and Tab. 2, which present long-term and short-term forecasting results respectively. The best results are highlighted in red and the second-best are underlined. The lower the MSE/MAE, the more accurate the forecast. AMD stands out with the best performance in 50 cases and the second best in 27 cases out of the overall 80 cases. Compared with other baselines, AMD performs well on both high-dimensional and low-dimensional datasets. It is worth noting that PatchTST does not perform well on the PEMS dataset, possibly because the patching design leads to the neglect of highly fluctuating temporal patterns. In contrast, AMD leverages information from multi-scale temporal patterns. Furthermore, the performance of Crossformer is unsatisfactory, possibly because it introduces unnecessary noise by exploring cross-variable dependencies. AMD skillfully reveals the intricate dependencies existing among time steps across various variables. Additionally, DLinear and MTS-mixers perform poorly on high-dimensional datasets, whereas AMD can handle them. Moreover, as shown in Fig. 3(a), in most cases, the forecasting performance benefits from the increase of input length, which is also observed in the majority of other datasets.
Table 1: Long-term forecasting task. All the results are averaged from 4 different prediction lengths $T\in\{96,192,336,720\}$. To the best for a fairer comparison for all baselines, the input sequence length $L$ is searched among $\{96,192,336,512,672,720\}$. See Appendix D.1 for the full results.
Models AMD PatchTST Crossformer FEDformer TimesNet MICN DLinear MTS-Mixers
[6]
[9]
[8]
[21]
[20]
[13]
[16]
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Weather
0.223 0.262 0.226 0.264 0.230 0.290 0.310 0.357 0.259 0.287 0.242 0.298 0.240 0.300 0.235 0.272
ETTh1
0.407 0.424 0.413 0.431 0.441 0.465 0.428 0.454 0.458 0.450 0.433 0.462 0.423 0.437 0.430 0.436
ETTh2
0.350 0.392 0.330 0.379 0.835 0.676 0.388 0.434 0.414 0.427 0.385 0.430 0.431 0.447 0.386 0.413
ETTm1
0.347 0.375 0.351 0.381 0.431 0.443 0.382 0.422 0.400 0.406 0.383 0.406 0.357 0.379 0.370 0.395
ETTm2
0.253 0.315 0.255 0.315 0.632 0.578 0.292 0.343 0.291 0.333 0.277 0.336 0.267 0.332 0.277 0.325
ECL
0.159 0.254 0.159 0.253 0.293 0.351 0.207 0.321 0.192 0.295 0.182 0.297 0.177 0.274 0.173 0.272
Exchange
0.328 0.387 0.387 0.419 0.701 0.633 0.478 0.478 0.416 0.443 0.315 0.404 0.297 0.378 0.373 0.407
Traffic
0.393 0.272 0.391 0.264 0.535 0.300 0.604 0.372 0.620 0.336 0.535 0.312 0.434 0.295 0.494 0.354
Solar-Energy
0.192 0.240 0.256 0.298 0.204 0.248 0.243 0.350 0.244 0.334 0.213 0.266 0.329 0.400 0.315 0.363
Table 2: Short-term forecasting task. The input sequence length $L$ is 96, while the prediction length $T$ is 12 for all baselines.
Models AMD PatchTST Crossformer FEDformer TimesNet TiDE DLinear MTS-Mixers
[6]
[9]
[8]
[21]
[12]
[13]
[16]
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
@@ Results.
0.093 0.206 0.168 0.232 0.165 0.214 0.173 0.273 0.112 0.212 0.227 0.343 0.154 0.276 0.186 0.286
5.2 Ablation Study and Analysis
Figure 3: (a) The performance on Weather dataset of AMD improves as the input length increases, indicating our model can effectively extract useful information from longer history and capture long-term multi-scale dependency. (b) Cross-channel dependencies may lead to deviations from the original distribution.
Observation 1: AMS learn from different temporal patterns.
As shown in Tab. 3, we demonstrate the performance improvement attributed to AMD is not solely due to the enlarged model size, but rather the integration of temporal pattern information, we devised the following experiments: (1) The permutation-invariant nature of the self-attention mechanism leads to temporal information loss, whereas the MLP-based model inherently maintains the sequential order of data. To demonstrate the utilization of sequential information, we replaced the temporal pattern embedding with random orders as RandomOrder. (2) To demonstrate the effectiveness of AMS aggregation, we replaced the TP-Selector with the average weighting as AverageWeight, treating different temporal patterns equally.
Consistent with our observation, RandomOrder and AverageWeight produce losses higher than AMD in most cases. Compared to self-attention, whose permutation-invariant nature leads to the loss of sequential information, AMS makes better use of temporal relations. Compared to simple averaging, AMS adaptively assigns corresponding weights to different temporal patterns, resulting in more accurate predictions.
On top of that, AMD benefits from improved interpretability. We plotted selector weights as shown in Fig. 4(a). Before and after the time step T, the temporal variations are respectively dominated by TP16 and TP5. Before T, the predicted data resembles the trend of TP16, both exhibiting a downward fluctuation. However, after T, the predicted data resembles TP5, which suddenly shows a significant increase. AMD recognizes this changing dominant role over time and adaptively assigns them higher weights.
Observation 2: Cross-channel dependencies are not always effective.
To prove cross-channel dependencies often introduce unwanted noise, we conducted another ablation study on the Weather dataset as shown in Tab. 3. We introduce cross-channel dependencies by adjusting the scaling rate. Specifically, we set the parameter $\beta$ in Eq. 7 to 0.5 and 1.0, respectively. In addition, we conduct experiments by removing components (w/o) DDI. From the results, it can be seen that the introduction of cross-channel dependencies do not enhance the prediction accuracy, and this was consistent across other datasets as well.
To explain this phenomenon, we visualize the learned dependencies as shown in Fig. 3(b). Compared to the temporal dependency, especially when the target variable is not correlated with other covariates, cross-channel dependencies tend to smooth out the variability in the target variable, causing its distribution to deviate from what would be expected based solely on its own past values.
Table 3: Component ablation of AMD on Weather and ECL.
@@ 5.2 Ablation Study and Analysis
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Weather 96 0.145 0.198 0.152 0.209 0.149 0.205 0.154 0.208 0.146 0.198 0.147 0.200 0.160 0.211
192 0.190 0.240 0.194 0.245 0.194 0.246 0.197 0.248 0.194 0.245 0.195 0.246 0.209 0.258
336 0.242 0.280 0.245 0.286 0.249 0.288 0.249 0.289 0.249 0.287 0.248 0.286 0.278 0.309
720 0.315 0.332 0.321 0.338 0.325 0.340 0.323 0.340 0.318 0.335 0.320 0.336 0.364 0.366
ECL 96 0.129 0.225 0.133 0.230 0.135 0.234 0.132 0.231 0.140 0.242 0.147 0.248 0.139 0.238
192 0.148 0.242 0.155 0.248 0.155 0.251 0.151 0.247 0.169 0.268 0.175 0.273 0.156 0.456
336 0.164 0.258 0.171 0.264 0.169 0.263 0.167 0.266 0.174 0.274 0.183 0.285 0.177 0.271
720 0.195 0.289 0.202 0.294 0.202 0.295 0.200 0.294 0.207 0.298 0.215 0.308 0.208 0.299
Observation 3: TP-Selector helps not only generalization but also efficiency.
We thoroughly compare the training time and memory usage of various baselines on the ETTh1 dataset, using the official model configurations and the same batch size. The results, shown in Fig. 4(b), indicate that AMD demonstrates superior efficiency with a relatively small number of parameters.

### v2

@@ Introduction
where $k$ is the number of dominant temporal patterns, $\psi\in\mathbb{N}(0,1)$ is standard Gaussian noise, and $\mathbf{W}_{\text{noise}}\in\mathbb{R}^{m\times m}$ is a learnable weight controlling the noisy values.
Models
AMD (Ours) TimeMixer PatchTST iTransformer Crossformer FEDformer TimesNet MICN DLinear MTS-Mixers
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Weather 96 0.145 0.197 0.147 0.197 0.149 0.198 0.174 0.214 0.153 0.217 0.238 0.314 0.172 0.220 0.161 0.226 0.152 0.237 0.156 0.206
192 0.187 0.238 0.189 0.239 0.194 0.241 0.221 0.254 0.197 0.269 0.275 0.329 0.219 0.261 0.220 0.283 0.220 0.282 0.199 0.248
336 0.240 0.280 0.241 0.280 0.245 0.282 0.278 0.296 0.252 0.311 0.339 0.377 0.280 0.306 0.275 0.328 0.265 0.319 0.249 0.291
720 0.315 0.330 0.310 0.330 0.314 0.334 0.358 0.349 0.318 0.363 0.389 0.409 0.365 0.359 0.311 0.356 0.323 0.362 0.336 0.343
ETTh1 96 0.369 0.397 0.361 0.390 0.370 0.399 0.386 0.405 0.386 0.429 0.376 0.415 0.384 0.402 0.396 0.427 0.375 0.399 0.372 0.395
192 0.401 0.416 0.409 0.414 0.413 0.421 0.441 0.436 0.416 0.444 0.423 0.446 0.457 0.436 0.430 0.453 0.405 0.416 0.416 0.426
336 0.418 0.427 0.430 0.429 0.422 0.436 0.487 0.458 0.440 0.461 0.444 0.462 0.491 0.469 0.433 0.458 0.439 0.443 0.455 0.449
720 0.439 0.454 0.445 0.460 0.447 0.466 0.503 0.491 0.519 0.524 0.469 0.492 0.521 0.500 0.474 0.508 0.472 0.490 0.475 0.472
ETTh2 96 0.274 0.337 0.271 0.330 0.274 0.337 0.297 0.349 0.628 0.563 0.332 0.374 0.340 0.374 0.289 0.357 0.289 0.353 0.307 0.354
192 0.351 0.383 0.317 0.402 0.339 0.379 0.380 0.400 0.703 0.624 0.407 0.446 0.402 0.414 0.409 0.438 0.383 0.418 0.374 0.399
336 0.375 0.411 0.332 0.396 0.329 0.380 0.428 0.432 0.827 0.675 0.400 0.447 0.452 0.452 0.417 0.452 0.448 0.465 0.398 0.432
720 0.402 0.438 0.342 0.408 0.379 0.422 0.427 0.445 1.181 0.840 0.412 0.469 0.462 0.468 0.426 0.473 0.605 0.551 0.463 0.465
ETTm1 96 0.284 0.339 0.291 0.340 0.290 0.342 0.334 0.368 0.316 0.373 0.326 0.390 0.338 0.375 0.314 0.360 0.299 0.343 0.314 0.358
192 0.322 0.362 0.327 0.365 0.332 0.369 0.377 0.391 0.377 0.411 0.365 0.415 0.371 0.387 0.359 0.387 0.335 0.365 0.354 0.386
336 0.360 0.380 0.360 0.381 0.366 0.392 0.426 0.420 0.431 0.442 0.391 0.425 0.410 0.411 0.398 0.413 0.369 0.386 0.384 0.405
720 0.421 0.416 0.415 0.417 0.416 0.420 0.491 0.459 0.600 0.547 0.446 0.458 0.478 0.450 0.459 0.464 0.425 0.421 0.427 0.432
ETTm2 96 0.167 0.258 0.164 0.254 0.166 0.256 0.180 0.264 0.421 0.461 0.180 0.271 0.187 0.267 0.178 0.273 0.167 0.260 0.177 0.259
192 0.221 0.294 0.223 0.295 0.223 0.296 0.250 0.309 0.503 0.519 0.252 0.318 0.249 0.309 0.245 0.316 0.224 0.303 0.241 0.303
336 0.270 0.327 0.279 0.330 0.274 0.329 0.311 0.348 0.611 0.580 0.324 0.364 0.321 0.351 0.295 0.350 0.281 0.342 0.297 0.338
720 0.356 0.382 0.359 0.383 0.362 0.385 0.412 0.407 0.996 0.750 0.410 0.420 0.497 0.403 0.389 0.409 0.397 0.421 0.396 0.398
ECL 96 0.129 0.224 0.129 0.224 0.129 0.222 0.148 0.240 0.187 0.283 0.186 0.302 0.168 0.272 0.159 0.267 0.153 0.237 0.141 0.243
192 0.147 0.238 0.140 0.220 0.147 0.240 0.162 0.253 0.258 0.330 0.197 0.311 0.184 0.289 0.168 0.279 0.152 0.249 0.163 0.261
336 0.160 0.253 0.161 0.255 0.163 0.259 0.178 0.269 0.323 0.369 0.213 0.328 0.198 0.300 0.196 0.308 0.169 0.267 0.176 0.277
720 0.193 0.286 0.194 0.287 0.197 0.290 0.225 0.317 0.404 0.423 0.233 0.344 0.220 0.320 0.203 0.312 0.233 0.344 0.212 0.308
Exchange 96 0.083 0.201 0.086 0.204 0.093 0.214 0.086 0.206 0.186 0.346 0.136 0.276 0.107 0.234 0.102 0.235 0.081 0.203 0.083 0.201
192 0.171 0.293 0.176 0.298 0.192 0.312 0.177 0.299 0.467 0.522 0.256 0.369 0.226 0.344 0.172 0.316 0.157 0.293 0.174 0.296
336 0.309 0.402 0.345 0.426 0.350 0.432 0.331 0.417 0.783 0.721 0.426 0.464 0.367 0.448 0.272 0.407 0.305 0.414 0.336 0.417
720 0.750 0.652 0.848 0.692 0.911 0.716 0.847 0.691 1.367 0.943 1.090 0.800 0.964 0.746 0.714 0.658 0.643 0.601 0.900 0.715
Traffic 96 0.366 0.259 0.360 0.249 0.360 0.249 0.395 0.268 0.512 0.290 0.576 0.359 0.593 0.321 0.508 0.301 0.410 0.282 0.462 0.332
192 0.381 0.265 0.379 0.256 0.375 0.350 0.417 0.276 0.523 0.297 0.610 0.380 0.617 0.336 0.536 0.315 0.423 0.287 0.488 0.354
336 0.397 0.269 0.385 0.270 0.392 0.264 0.433 0.283 0.530 0.300 0.608 0.375 0.629 0.336 0.525 0.310 0.436 0.296 0.498 0.360
720 0.429 0.292 0.430 0.281 0.432 0.286 0.467 0.302 0.573 0.313 0.621 0.375 0.640 0.350 0.571 0.323 0.466 0.315 0.529 0.370
Solar 96 0.175 0.228 0.167 0.220 0.224 0.278 0.203 0.237 0.181 0.240 0.201 0.304 0.219 0.314 0.188 0.252 0.289 0.337 0.284 0.325
192 0.186 0.235 0.187 0.249 0.253 0.298 0.233 0.261 0.196 0.252 0.237 0.337 0.231 0.322 0.215 0.280 0.319 0.397 0.307 0.362
336 0.200 0.246 0.200 0.258 0.273 0.306 0.248 0.273 0.216 0.243 0.254 0.362 0.246 0.337 0.222 0.267 0.352 0.415 0.333 0.384
720 0.203 0.249 0.215 0.250 0.272 0.308 0.249 0.275 0.220 0.256 0.280 0.397 0.280 0.363 0.226 0.264 0.356 0.412 0.335 0.383
Table 1: Long-term forecasting task. All the results are averaged from 4 different prediction lengths $T\in\{96,192,336,720\}$. To the best for a fairer comparison for all baselines, the input sequence length $L$ is searched among $\{96,192,336,512,672,720\}$.
@@
TP-Projection takes the embedding $\mathbf{v}$ from the output $\mathbf{V}$ of DDI as input. Each predictor consists of multiple feedforward layers. These outputs are then multiplied by the $\mathbf{S}$ and summed to yield the prediction result $\mathbf{\hat{y}}$ for one channel. The final results $\mathbf{\hat{Y}}$ are composed of the outputs $\mathbf{\hat{y}}$ from each channel:
@@
Compared to sparse MoE (Li et al. 2023a), we adopt dense MoE in TP-Projection with two considerations. Firstly, each temporal pattern contributes to the prediction result. Secondly, this approach can help mitigate issues such as load unbalancing and embedding omissions that are prevalent in sparse MoE architectures. Our TopK method can be formalized with:
@@
where $\|\mathbf{\Theta}\|_{2}$ is the L2-norm, $\lambda_{1,2}$ are hyper-parameters.
Experiments
Main Results
@@
We conduct experiments on seven real-world datasets, including Weather, ETT (ETTh1, ETTh2, ETTm1, ETTm2), ECL, Exchange, Traffic and Solar Energy for long-term forecasting and PEMS (PEMS03, PEMS04, PEMS07, PEMS08) for short-term forecasting.
@@
We carefully select some representative models to serve as baselines in the field of time series forecasting, including (1) MLP-based models: TimeMixer (Wang et al. 2024), TiDE (Das et al. 2023), MTS-Mixers (Li et al. 2023b), RLinear (Li et al. 2024), and DLinear (Zeng et al. 2023); (2) Transformer-based models: PatchTST (Nie et al. 2023), iTransformer (Liu et al. 2024a), Crossformer (Zhang and Yan 2023), and FEDformer (Zhou et al. 2022b); (3) CNN-based models: TimesNet (Wu et al. 2023), and MICN (Wang et al. 2023).
@@
To ensure fair comparisons, for long-term forecasting, we rerun all baselines with different input lengths $L$ and choose the best results to avoid underestimating the baselines. For short-term forecasting, the input length is 96. We select two common metrics in time series forecasting: Mean Absolute Error (MAE) and Mean Squared Error (MSE). All experiments are conducted using PyTorch on an NVIDIA V100 32GB GPU and are repeated five times for consistency.
Models
AMD PatchTST Crossformer FEDformer TimesNet TiDE DLinear MTS-Mixers RLinear
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
@@
0.093 0.206 0.168 0.232 0.165 0.214 0.173 0.273 0.112 0.212 0.227 0.343 0.154 0.276 0.186 0.286 0.133 0.247
Table 2: Short-term forecasting task. The input sequence length $L$ is 96 and the prediction length $T$ is 12.
@@
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Weather 96 0.145 0.198 0.150 0.205 0.152 0.209 0.149 0.205 0.154 0.208 0.146 0.198 0.147 0.200 0.160 0.211 0.149 0.204 0.150 0.206
192 0.187 0.238 0.193 0.245 0.194 0.245 0.194 0.246 0.197 0.248 0.194 0.245 0.195 0.246 0.209 0.258 0.192 0.243 0.199 0.248
336 0.240 0.280 0.244 0.284 0.245 0.286 0.249 0.288 0.249 0.289 0.249 0.287 0.248 0.286 0.278 0.309 0.246 0.285 0.248 0.287
720 0.315 0.330 0.315 0.334 0.321 0.338 0.325 0.340 0.323 0.340 0.318 0.335 0.320 0.336 0.364 0.366 0.319 0.336 0.320 0.339
ECL 96 0.129 0.224 0.132 0.228 0.133 0.230 0.135 0.234 0.132 0.231 0.140 0.242 0.147 0.248 0.139 0.238 0.135 0.231 0.137 0.234
192 0.147 0.238 0.151 0.244 0.155 0.248 0.155 0.251 0.151 0.247 0.169 0.268 0.175 0.273 0.156 0.456 0.153 0.247 0.157 0.251
336 0.160 0.253 0.165 0.259 0.171 0.264 0.169 0.263 0.167 0.266 0.174 0.274 0.183 0.285 0.177 0.271 0.169 0.265 0.171 0.267
720 0.193 0.286 0.199 0.293 0.202 0.294 0.202 0.295 0.200 0.294 0.207 0.298 0.215 0.308 0.208 0.299 0.201 0.296 0.203 0.298
Table 3: Component ablation of AMD on Weather and ECL.
Results.
Comprehensive forecasting results are shown in Tab. 1 and Tab. 2, which present long-term and short-term forecasting results respectively. The best results are highlighted in red and the second-best are underlined. The lower the MSE/MAE, the more accurate the forecast. AMD stands out with the best performance in 50 cases and the second best in 27 cases out of the overall 80 cases. Compared with other baselines, AMD performs well on both high-dimensional and low-dimensional datasets. It is worth noting that PatchTST does not perform well on the PEMS datasets for the patching design leads to the neglect of highly fluctuating temporal patterns. In contrast, AMD leverages information from multi-scale temporal patterns. Furthermore, the performance of Crossformer is unsatisfactory for it introduces unnecessary noise by exploring cross-variable dependencies. AMD skillfully reveals the intricate dependencies existing among time steps across various variables. Additionally, DLinear and MTS-mixers perform poorly on high-dimensional datasets, whereas AMD can handle them.
Model Analysis
Ablation Study.
To verify the effectiveness of each component of AMD, we perform detailed ablation of each possible design on Weather and ECL datasets. As shown in Tab. 3, we have the following observations.
For MDM, we conduct experiments by removing MDM (w/o MDM). The increased error suggests that relying on a single scale is insufficient for accurate prediction.
For DDI, recent studies (Nie et al. 2023; Zhang and Yan 2023) show that both channel independence and cross-channel dependencies strategies can achieve SOTA accuracy in specific tasks. Typically, datasets with strong cross-channel correlations perform better using cross-channel dependencies strategies. However, for datasets with weak inter-channel correlations, cross-channel dependencies often introduce unwanted noise. To prove this, we introduce cross-channel dependencies by adjusting the scaling rate. Specifically, we set the parameter $\beta$ in Eq. 7 to 0.5 and 1.0, respectively. In addition, we conduct experiments by removing DDI (w/o DDI). From the results, it can be seen that the introduction of cross-channel dependencies do not always enhance the prediction accuracy, and this was consistent across other datasets as well. Scaling rate $\beta$ can balance the emphasis on temporal dependencies and cross-channel dependencies in multivariate time series data and boost the model performance.
Figure 3: AMD guides the prediction by assigning greater weight to the dominant time pattern.
For AMS, we demonstrate the performance improvement is not solely due to the enlarged model size, but rather the integration of temporal pattern information, we devised the following experiments: To demonstrate the utilization of sequential information, we replaced the temporal pattern embedding with random orders as RandomOrder. To demonstrate the effectiveness of AMS aggregation, we replaced the TP-Selector with the average weighting as AverageWeight, treating different temporal patterns equally. We also remove AMS (w/o AMS) and replace it with a single linear predictor. RandomOrder, AverageWeight and w/o AMS all cause a performance decline. This illustrates that, compared to self-attention whose permutation-invariant nature results in the loss of sequential information, AMS inherently maintains the sequential order and makes better use of temporal relationships by identifying dominant patterns that change over time. Compared to simple averaging like timemixer (Wang et al. 2024), AMS adaptively assigns corresponding weights to different temporal patterns, resulting in more accurate predictions.
Model Interpretablilty.
To provide an intuitive understanding of AMS, we plotted selector weights in Fig. 3. Compared to a single linear predictor, AMS offers stronger generalization capabilities and improved interpretability. Before and after the time step T, the temporal variations are respectively dominated by TP16 and TP5. Before T, the predicted data resembles the trend of TP16, both exhibiting a downward fluctuation. However, after T, the predicted data resembles TP5, which suddenly shows a significant increase. AMD recognizes this changing dominant role over time and adaptively assigns them higher weights.
Figure 4: Cross-channel dependencies may lead to deviations from the original distribution.
To offer a clear conceptualization of the balance of emphasis on temporal dependencies and cross-channel dependencies mentioned above, we visualize the learned dependencies as shown in Fig. 4. Compared to the temporal dependencies, especially when the target variable is not correlated with other covariates, cross-channel dependencies tend to smooth out the variability in the target variable, causing its distribution to deviate from what would be expected based solely on its own past values.